OpenStudy (mmend98):

The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate? (5 points) f(x) = 4(4)x; growth rate 400% f(x) = 4(4)3x, growth rate 4% f(x) = 4(1.59)3x; growth rate 59% f(x) = 4(1.59)x; growth rate 4%

OpenStudy (thomas5267):

What is f(x) equal to? What is 4(4)x?

OpenStudy (mmend98):

I think it would be B.

OpenStudy (mmend98):

they are exponential as well the answers sorry for the typo

ganeshie8 (ganeshie8):

rearrange the exponent as \(\frac{3x}{3}\) and simplify : \[\large{\begin{align}f(x)=4(4)^x &= 4(4)^{\frac{3x}{3}}\\~\\ &= 4(4^{\frac{1}{3}})^{3x}\\~\\ &= ? \end{align}}\]

OpenStudy (mmend98):

ok

ganeshie8 (ganeshie8):

use ur calculator to work \(\large 4^{\frac{1}{3}}\)

OpenStudy (mmend98):

16.3

ganeshie8 (ganeshie8):

nope, try again

ganeshie8 (ganeshie8):

wolf says \(\approx 1.59\) http://www.wolframalpha.com/input/?i=4%5E%281%2F3%29

OpenStudy (mmend98):

ah

OpenStudy (mmend98):

So the answer would be D?

OpenStudy (mmend98):

or C.

OpenStudy (mmend98):

c would be the answer

ganeshie8 (ganeshie8):

right, one of them how do we pin down which one it is

OpenStudy (mmend98):

it is c?

ganeshie8 (ganeshie8):

how do you know

OpenStudy (mmend98):

because it is the same but manipulatd as the original equation

ganeshie8 (ganeshie8):

rearrange the exponent as \(\frac{3x}{3}\) and simplify : \[\large{\begin{align}f(x)=4(4)^x &= 4(4)^{\frac{3x}{3}}\\~\\ &= 4(4^{\frac{1}{3}})^{3x}\\~\\ &= 4(1.\color{red}{59})^{3x}\\~\\ \end{align}}\] that red part gives you the growth rate

OpenStudy (mmend98):

oh ok.

ganeshie8 (ganeshie8):

the fly population grows at a rate of \(59\%\) per quarter

OpenStudy (mmend98):

we know c is the anser because 59 percent is the growht rate

OpenStudy (mmend98):

?

ganeshie8 (ganeshie8):

yes ?

OpenStudy (mmend98):

yes

OpenStudy (irishboy123):

.